# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v2

@@ 3.3 Reward Model Mis-specification
Under Condition 3, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{\Pi}^{\star}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Observation under token-level parameterization. To assess the practicality of Condition 3 for auto-regressive language models, we specialize our general bandit model to the token-level parameterization. In this setting, the optimal policy admits the closed-form characterization of Rafailov et al. (2024), which we restate with an explicit separation between $\pi_{{\mathsf{ref}}}$ and the $q^{\star}$ function (see Section B.11 for details):
@@
[…]
This observation shows that while the reward model in RLHF only needs to approximate $r^{\star}$, the policy model in DPO must capture the token-level $q^{\star}$ function, which recursively entangles the reward signal with the base model $\pi_{{\mathsf{ref}}}$. As a result, the policy model faces a substantially more demanding learning objective, making it more prone to mis-specification and suggesting that the “weak reward, strong policy model” regime may be less common in practice.
3.4 Double Model Mis-specification
We now consider the most challenging setting, where neither the ground-truth reward function nor the optimal policy is realizable by their respective model classes.
@@ 3.4 Double Model Mis-specification
Analysis of the isomorphic case. Condition 5 indicates the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies. This structure allows us to directly compare RLHF and DPO when both operate under the same representational constraints, and to investigate whether bypassing reward modeling, as in DPO, provides any advantage. In RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling. Therefore, RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO. Proofs deferred to Sections B.5 and B.9.
@@
On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections B.6 and B.7.
@@
Though there is no consistent performance gap between RLHF and DPO in certain settings, revisiting the framework can reveal a structural parallel: RLHF can yield the best policy given the learned reward model $r_{\textup{RLHF}}$, and the DPO policy is directly the optimal one given the surrogate reward model $\hat{r}_{\textup{DPO}}$. And online DPO serves to enhance the quality of $\hat{r}_{\textup{DPO}}$ (Xiong et al., 2024). Formally,
@@
This result implies a general principle: the performance gap is reflected in the quality gap between the (surrogate) reward models: $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$. Better reward learning yields higher expected value.
As revealed in Appendix C of Ouyang et al. (2022) and Section 3.3 of Swamy et al. (2025), it is uncommon to deploy a reward model with a larger scale than the policy model. And thus to ensure practical relevance, we focus on the regime $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
As an answer to this question, we note that in the context of preference learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference, derived by simple calculations:
$V_{r_{\phi}}^{\pi_{\theta^{\star}(r_{\phi})}}\textit{ can be measured by }-\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}\;,$ (3)
where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$ and we omit prompts for simplicity. Detailed calculations and further discussions deferred to Appendix A. Using this metric, we can further establish a separation.
Concluding remarks. Although we adopt relatively simple techniques, these results can provide valuable insights for the fundamental differences between RLHF and DPO. In the next section, we demonstrate that these insights extend naturally to more practical and realistic scenarios.
@@
With limited preference data, we are not able to directly compute exact solutions, and thus obtain weaker reward models and policy models due to estimation error. This scenario can be viewed as inducing an implicit model mis-specification, whose effects have been widely discussed in Condition 4 of Section 3. And since we can only lose information in reward learning (Swamy et al., 2025), Equation 2 still holds asymptotically with on-policy sampling. Thus by assuming $F\subseteq F_{\Pi}$, we only need to compare the reward model quality measure shown in Equation 3. We adopt an empirical proxy for this notion, data-induced semi-norm (details in Definition 2 in Section B.10, see also Zhu et al. (2023)): $\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\right]^{2}$, where $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n}$ is an empirical preference dataset and we omit prompts from now on.
@@
Difference in token-level linear parameterization. In this section, to rigorously establish a separation, we focus on a specific token-level linear parameterization, which is a special case of the general bandit model; therefore, previous results continue to hold. The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value. Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta_{r}}(y)=\beta\sum_{t=0}^{n}{\theta_{r,t}}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and n=|y|1n=|y|-1. While for the policy model, one needs to go through the softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al. (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
$\pi_{\theta_{p}}(y)=\prod_{t=0}^{n}\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{n}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
@@
Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{n}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$, then the optimal solution for the reward model is $\theta_{r,t}^{\star}=\theta^{\star}_{t}$. And recall Equation 1, the optimal solution for the policy model is:
@@
$\pi_{\theta^{\star}_{p,t}}(y_{t}|y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;.$
Benefiting from the token-level $q^{\star}$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025). However, simultaneously, learning the $q^{\star}$ function sacrifices statistical efficiency due to the need to model the complicated structure. Next, we will present a concrete example to illustrate the statistical gap between pure reward learning and surrogate reward learning.
@@
Dual-token sparse prediction (DTSP) task.   The policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
We let $\theta_{r}^{\star}$ denote the optimal solution for pure reward learning, and $\theta_{p}^{\star}$ the optimal solution for surrogate reward learning. Note that for the second token, $\theta_{r}^{\star}$ and $\theta_{p}^{\star}$ share the same optimal solution:
$(\theta_{r,1}^{\star})^{\top}\psi(y,\omega)=e_{1}^{\top}\psi(y,\omega)+C_{1}\;,\;(\theta_{p,1}^{\star})^{\top}\psi(y,\omega)=e_{1}^{\top}\psi(y,\omega)+C_{2}\;,$
where $C_{1},C_{2}\in\mathbb{R}$ are offsets. And for the first token $y$, there is a distinction:
$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
where $\textbf{r}_{\textup{sparse}}$ gets entangled with $\pi_{{\mathsf{ref}}}$ in $\theta_{p,0}^{\star}$. Note that if $\log\mathbb{E}_{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}\exp(\psi(y,\omega)_{1})$ can be mapped to certain complicated non-linear function of $\psi(y)$, then the policy model is mis-specified while the reward model is not, as in Condition 2. And even without explicit model mis-specification, we can establish a separation in reward model qualities due to statistical efficiency differences.
Theorem 10 (Informal).
@@
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r}}(y_{w}^{(i)})-r_{\hat{\theta}_{r}}(y_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
@@
By fixing the optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to a single-token prediction problem, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense. Leveraging the results of Yao et al. (2025) then yields the separation. Formal statement and detailed proof deferred to Section B.10.
Concluding remarks. This section shows that the estimation error can also induce an implicit model mis-specification. From the perspective of sparse recovery, we can see that the DPO could suffer from severe statistical inefficiency compared with pure reward learning, even with the same model scale. Although our task construction is specific, it reveals a general phenomenon: DPO can distort the intrinsic structure of the true reward function. Given the insight that real-world rewards are often sparse and simple (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the surrogate reward model. This further explains why two-stage RLHF is empirically observed to outperform DPO (Ivison et al., 2024; Xu et al., 2024b).

### v3

@@ 3 Exact Optimization: Fine-grained Performance Gap Induced by Model Mis-specification
Under Condition 3, $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}=V^{\Pi}_{r^{\star}}$, and there exists an environment s.t. $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$.
Observation under token-level parameterization. To assess the practicality of Conditions 2 and 3 for auto-regressive language models, we can specialize our general policy class to the token-level parameterization. In this setting, the optimal policy admits the closed-form characterization of Rafailov et al. (2024), which we restate with an explicit separation between $\pi_{{\mathsf{ref}}}$ and the $q^{\star}$ function (see Appendix B.10 for details):
@@
$q^{\star}(y_{t}|x,y_{0\ldots t-1})=\begin{cases}\beta\log\mathbb{E}_{s\sim\pi_{\textup{ref}}(\cdot|x,y_{0\ldots t})}\exp\!\big(q^{\star}(s|x,y_{0\ldots t})/\beta\big)\;\;&\text{if $y_{t}$ is not terminal,}\\ r^{\star}(x,y_{0\ldots t})\;\;\begin{subarray}{c}\end{subarray}&\text{if $y_{t}$ is terminal}.\end{cases}$
Here $\mathcal{V}$ is the vocabulary, and $s\in\mathcal{V}$ is the token. This observation shows that while the reward model in RLHF only needs to approximate $r^{\star}$, the policy model in DPO must capture the token-level $q^{\star}$ function, which recursively entangles the reward signal with the base model $\pi_{{\mathsf{ref}}}$. As a result, the policy model faces a substantially more demanding learning objective, making it more prone to mis-specification than the reward model of the same scale. Therefore, in practice, people sometimes deploy small reward model and can still obtain acceptable performance, while the deployed policy model cannot be too small.
Double Model Mis-specification. We now consider the most challenging setting, where neither the ground-truth reward function nor the optimal policy is realizable by their respective model classes.
@@
Analysis of the isomorphic case. Condition 5 indicates the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies. This structure allows us to directly compare RLHF and DPO when both operate under the same representational constraints, and to investigate whether bypassing reward modeling, as in DPO, provides any advantage. In RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling. Therefore, RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO. Proofs deferred to Appendix B.5 and  B.9.
@@
On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Appendix B.6 and  B.7.
@@
Understanding the role of reward quality. Though there is no consistent performance gap between RLHF and DPO in certain settings, revisiting the framework can reveal a structural parallel: RLHF can yield the best policy given the learned reward model $r_{\textup{RLHF}}$, and the DPO policy is directly the optimal one given the surrogate reward model $\hat{r}_{\textup{DPO}}$. And online DPO/RLHF (Xiong et al., 2024; Li et al., 2025) serves to enhance the quality of $\hat{r}_{\textup{DPO}}$. Formally,
@@
This result implies a general principle holding for all conditions in this section: the performance gap is reflected in the quality gap between the (surrogate) reward models: $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$. Better reward learning yields higher expected value. We will look into reward model qualities in next section, and defer an algorithmic discussion to Appendix A.
Concluding remarks. Although we adopt relatively simple techniques, these results can provide valuable insights for the fundamental differences between RLHF and DPO. Later we will demonstrate that these insights extend naturally to more practical and realistic scenarios.
@@
With limited preference data, we are not able to directly compute exact solutions, and thus obtain weaker reward models and policy models due to estimation error. This scenario can be viewed as inducing an implicit model mis-specification, whose effects have been widely discussed in Section 3. And because we can only lose information in reward learning (Swamy et al., 2025), Equation 2 still holds asymptotically with on-policy sampling. So our main focus will be to compare the qualities of reward model and DPO model, and we omit prompts from now on.
Notably, for linear bandit model studied in previous studies (Xiong et al., 2024; Zhu et al., 2023; Yao et al., 2025), where the ground truth reward is $r^{\star}(y)=\beta(\theta^{\star})^{\top}\psi(y)$, the reward model is $r_{\theta_{r}}(y)=\beta\theta^{\top}\psi(y)$, and the policy model is $\pi_{\theta_{p}}(y)\propto\pi_{{\mathsf{ref}}}(y)\exp(\theta^{\top}\psi(y))$, we have that $\theta_{r}$ and $\theta_{p}$ share the same optimal solution $\theta^{\star}$, and there is no difference in sample complexity. To rigorously establish a separation, we focus on the more realistic token-level linear parameterization, which is a special case of our general policy class; therefore, results in Section 3 continue to hold.
@@
Reward model. The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value. Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.:
$r_{\theta_{r}}(y)=\beta\sum_{t=0}^{|y|-1}\theta_{r,t}^{\top}\psi(y_{0\ldots t})\;,$
where $\theta_{r,t},\psi(y_{0\ldots t})\in\mathbb{R}^{d}$.
Policy model. While for the policy model, one needs to go through softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al. (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
$\quad\pi_{\theta_{p}}(y)=\prod_{t=0}^{|y|-1}\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{|y|-1}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s\in\mathcal{V}}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t-1},s))}\;,$
where $\theta_{p,t}\in\mathbb{R}^{d}$, $\mathcal{V}$ is the vocabulary, and the surrogate reward model is $\hat{r}_{\theta_{p}}(y)=\beta\sum_{t=0}^{|y|-1}\log\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})$.
@@
Let $\theta_{r}^{\star}$ be the optimal solution for pure reward learning, and $\theta_{p}^{\star}$ be the optimal solution for DPO. Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{|y|-1}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$. Then the optimal solution for the reward model is
$\theta_{r,t}^{\star}=\theta^{\star}_{t}+C\;,$ (3)
where $C$ is a constant offset, and recall Equation 1, the optimal solution for the policy model is:
@@
$\pi_{\theta^{\star}_{p,t}}(y_{t}|y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;,$
and thus
$\left(\theta_{p,t}^{\star}\right)^{\top}\psi(y_{0\ldots t})=q^{\star}(y_{t}|y_{0\ldots t-1})/\beta+C^{\prime}=\left(\theta_{t}^{\star}\right)^{\top}\psi(y_{0\ldots t})+v^{\star}(y_{t}|y_{0\ldots t-1})/\beta+C^{\prime\prime}\;,$ (4)
where $v^{\star}$ is the soft value function (see detailed derivation in Appendix B.10), and $C^{\prime},C^{\prime\prime}$ are constant offsets.
Curse of value function. Comparing Equation 3 and Equation 4, we can see that the main distinction lies in whether to learn the value function. While it is widely recognized that learning $q^{\star}$$v^{\star}$ function can provide fine-grained process information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025), this benefit comes with a significant statistical cost: estimating $q^{\star}/v^{\star}$ generally incurs a higher sample complexity because the induced structure is considerably richer. Notably, even under a linear assumption for $q^{\star}/v^{\star}$, the value function couples the reward with the base model, so some good properties of the underlying reward (e.g., sparsity) are not preserved.
Next, we will present a minimalist example to rigorously illustrate the statistical gap between RLHF and DPO. We adopt an empirical proxy, data-induced semi-norm (details in Definition 2 in Appendix B.11, see also Zhu et al. (2023)): $\frac{1}{n}\sum_{i=1}^{n}\Delta^{2}_{y_{w}^{(i)},y_{l}^{(i)}}(r^{\star}-r_{\phi})$, where $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n}$ is an empirical preference dataset.
@@
Dual-token sparse prediction (DTSP) task.   Let $\mathcal{V}$ be the vocabulary, and $\mathcal{Y}=\mathcal{V}^{2}$. The policy model is required to sequentially output two tokens $a,b$, and the ground-truth reward is: $r^{\star}(a,b)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\beta e_{1}^{\top}\psi(a,b)\;,$ where $a,b\in\mathcal{V}$, $\psi(a),\psi(a,b)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
Theorem 10 (Informal reward separation).
@@
$\frac{1}{n}\sum_{i=1}^{n}\Delta^{2}_{y_{w}^{(i)},y_{l}^{(i)}}(r^{\star}-r_{\hat{\theta}_{r}})=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
@@
$\frac{1}{n}\sum_{i=1}^{n}\Delta^{2}_{y_{w}^{(i)},y_{l}^{(i)}}(r^{\star}-\hat{r}_{\hat{\theta}_{p}})=\Omega\left(\frac{d}{n}\right)\;.$
@@
By fixing the optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to a single-token prediction problem, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense. Leveraging the results of Yao et al. (2025) then yields the separation. Formal statement and detailed proof deferred to Appendix B.11.
Theorem 11 (Informal sub-optimality separation).
Under token-level linear parameterization and mild assumptions, there exists an environment for DTSP task, s.t. we can have a separation on the sub-optimality of RLHF and DPO:
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ $=\tilde{\mathcal{O}}\left(\sqrt[4]{\frac{k\log d}{n}}\cdot\sqrt{\Lambda_{1}}\right)\;,$
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $=\Omega\left(\frac{d}{n}\cdot\Lambda_{2}\right)\;,$
where $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\hat{\theta}_{r}}}^{\pi}$, $\pi_{\textup{DPO}}=\pi_{\hat{\theta}_{p}}$, and $\Lambda_{1},\Lambda_{2}$ are geometric quantities of data. Formal statement and detailed proof deferred to Appendix B.11.5.
Concluding remarks. In this section, we can see that the DPO could suffer from severe statistical inefficiency compared with pure reward learning, even under the same model scale. Although our parameterization is specific, it reveals a general phenomenon: DPO can distort the intrinsic structure of the true reward function. For general model class beyond linear model, Equation 1 still holds. And hence, compared with reward model, the policy model always faces a more complex target, making it more vulnerable to model mis-specification and sample inefficiency. Additionally, to prevent policy model mis-specification (discussed in Section 3), $d_{P}$ is often required to be larger than $d_{R}$, which further increases sample complexity. In the belief that real-world rewards are often simple or sparse (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the DPO model. This further explains why two-stage RLHF empirically outperforms DPO in practice (Ivison et al., 2024; Xu et al., 2024b).